Alexa01:

A lethal recessive condition is responsible for the death of 1 out of every 15,000 babies in a population. A. What is the frequency of the recessive genotype? B. What is the frequency of the heterozygous genotype? C. What is the frequency of the homozygous dominant genotype? D. What is the frequency of the dominant allele? E. What is the frequency of the recessive allele? F. Calculate the percent of the population that is heterozygous.

Madiron:

ok what is the question asking?

Alexa01:

It is asking for the calculations for A-F

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 A lethal recessive condition is responsible for the death of 1 out of every 15,000 babies in a population. A. What is the frequency of the recessive genotype? - so i think from these given details you can answer just the question A. bc. to answer the next questions of B,...,F you need to know more details again A. - a lethal recessive condition is responsible for the death of 1 out of every 15,000 babies in a population. using the rule of 3 simple you can write 1 death out of 15000 babies x death out of 100 babies cross multiplie and calcule the x value x = 100*1/15000 = 100/15000 = 1/150 = 0,00666 = 0,6*10^-2 % \[x = \frac{ 100*1 }{ 15000 } = \frac{ 100 }{ 15000 } = \frac{ 1 }{ 150 } = 0,0066 = 0,6*10^{-2}%\] \(\color{#0cbb34}{\text{End of Quote}}\) x = 0,6 *10^(-2) %